The electrical resistance, R, of a wire varies directly as its length and inversely as the square of its diameter. If the resistance of a 100 meter wire having a diameter of 1/2 cm. is 0.75 ohms, find the resistance of a 300 meter wire with a diameter of 1/4 cm.
Answer by stanbon(48558)

R=k/d^2
Need to find "k".
0.75=k/(1/2)^2
0.75(1/4)=k
k=(3/16)
EQUATION:
R=(3/16)/d^2
If d=1/4 find R.
R=(3/16)/(1/4)^2
R=3 ohms
Cheers,
Stan H.